OpenStudy (anonymous):

what equation represents a line passing through the points 2,-2 and -5,12

OpenStudy (precal):

do you mean (2,-2) and (-5,12)?

OpenStudy (anonymous):

Use this... y- y_1= m(x-x_1) where m = ( y_2-y_1 )/( x_2- x_1) (x_1,y_1) and (x_2,y_2) are two points passing throught the line.

OpenStudy (anonymous):

Substitute and get your answer

OpenStudy (precal):